Find Equilibrium Level Income and Rate of Interest from the following information: C-40+0.8 Yd, 110-5r, T-50, G-50, Md-0.2y-5r, Ms-100. Also calculate the impact of increase in govt. expenditure from Rs. 50 to Rs. 80 on income and interest rate.


1
Expert's answer
2021-07-15T10:38:08-0400

"Y = C + I + G = 40 + 0.8(Y - 50) + 110 - 5r + 50 = 0.8Y + 160 - 5r,"

0.2Y = 160 - 5r,

Ms = Md, so:

0.2Y - 5r = 100,

0.2Y = 100 + 5r,

160 - 5r = 100 + 5r,

10r = 60,

r = 6%.

"Y = (100 + 5*6)\/0.2 = 650."